4. (22) East is to the right and west is to the lef. At time to, Alice is at the origin, has a velocity of 2.50 m/s to the west and an acceleration 0.300 m/s east. At time to, Bob is 50.0 m east of Alice, and he has a velocity of 0.500 m/s east and an acceleration 0.120 m/s east. Alices displacement between time to and time ti is equal to Bobs displacement during this same time interval. a. What is the value of t?Circle your answer. b. What is Bobs position at time ti? Circle your answer. (Note: If you cannot find ti, then assume a reasonable value for ti to find Bobs position.) To receive full credit, you must show your work. The work must include an appropriate sketch and must use appropriate symbols. Looking for easy explanation and show of work for each step.
Answer #1

displacement = v0 t + a t^2 / 2

for Alice:

r = -2.50t + 0.30 t^2 /2

r = - 2.50t + 0.15 t^2

For Bob:

r = 0.5t + 0.120t^2 /2

equating both displacements,

- 2.50t + 0.15 t^2 = 0.5t + 0.120t^2 /2

0.09 t^2 - 3t = 0

t (0.09t - 3) = 0

t = 0 Or 33.33 sec

(A) t1 = 33.33 sec

(B) r = xf - xi = 0.5t + 0.120t^2 /2

xf - 50 = 0.5(33.33) + (0.120 x 33.33^2)/2

xf = 133.33 m